Projection Admin 9/11
To do this evening/ was completed: Had a little bit of a hassle with getting Apple TV up and running in T1 this morning for the Found Footage staff screening. The PC is not exactly reliable (mouse/wifi/and mind you, Apple TV app won't work on any of our iMacs that aren't 15 OS or higher) and I'll be looking into ways to make this less so. We only started a couple minutes late. I wound up being able to download the film so as to curtail any wifi issues.
We were emailed a new VF file for the Dubbed OCAP version of Demon Slayer. I updated the playlist.
For tomorrow - please tech Demon Slayer and Tropic Thunder. Thanks!
Presentation Issues:
Preshow Clips/Playlist: Built the playlists for Downton Abbey and scheduled the rest of the week with it.
Built the playlist for Thief of Joy, but we'll still need to tech the feature.
Trailer Updates:
Special Events/ AV set up: Went through the entire shorts playlist for this evening, which I had to make a few adjustments to. I noticed there was a 'house lights down' cue in the opening pack before the preshow, which I thought was odd. There was an incorrect lens/masking cue placement as well. Everything has been updated. This should be an easy, breezy show. They will need all four wireless mics, which are good to go. 8 chairs, too, but check with Desmond just in case there's a no show for one of the filmmakers. For stage right spots, use the 601 setting. Normal setting for center spots is on the sheet as usual.
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Alec forwarded me the DCP of Mars Express, which I downloaded via Aspera. This has since been ingested.
No updates, yet, to the Gotham shorts that still need uploading/fixing. I'll be checking in soon with the person who's helping to organize them. We still have quite a bit of time.
Outgoing/Shipping (content): Friendly reminder that we are waiting to ship out Vicious Lips (Vampire Lovers) so don't move anything yet!
Print/video Inspection: Got my eyes on Downton.
FYI - NewFest content should (fingers crossed) be here by the 17th, and our techs to start the 24th (of September!).
Asset tracking doc updated/checked?: We now have tracking for the Godfather PIII, which will be coming in three cans total, and will hopefully be here soon. One of the updates shows a delivery set for tomorrow afternoon, which would be great timing!
Maintenance/cleaning: Changed the lamp in T4 this morning, everything's looking nice and bright in there. Old lamp stored off to the side in our cave of explosives.
Cue sheets updated/checked (every shift): Yes.
Marquee updated?: All set.
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Asset tracking doc updated/checked?: We now have tracking for the Godfather PIII, which will be coming in three cans total, and will hopefully be here soon. One of the updates shows a delivery set for tomorrow afternoon, which would be great timing!
Maintenance/cleaning: Changed the lamp in T4 this morning, everything's looking nice and bright in there. Old lamp stored off to the side in our cave of explosives.
Cue sheets updated/checked (every shift): Yes.
Marquee updated?: All set.
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment